# Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Polyp Segmentation

### v1

@@ Abstract
Automatic polyp segmentation plays a crucial role in the early diagnosis and treatment of colorectal cancer (CRC). However, existing methods heavily rely on fully supervised training, which requires a large amount of labeled data with time-consuming pixel-wise annotations. Moreover, accurately segmenting polyps poses challenges due to variations in shape, size, and location. To address these issues, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised polyp Segmentation (DEC-Seg) from colonoscopy images. First, we propose a Cross-level Feature Aggregation (CFA) module that integrates cross-level adjacent layers to enhance the feature representation ability across different resolutions. To address scale variation, we present a scale-enhanced consistency constraint, which ensures consistency in the segmentation maps generated from the same input image at different scales. This constraint helps handle variations in polyp sizes and improves the robustness of the model. Additionally, we design a scale-aware perturbation consistency scheme to enhance the robustness of the mean teacher model. Furthermore, we propose a cross-generative consistency scheme, in which the original and perturbed images can be reconstructed using cross-segmentation maps. This consistency constraint allows us to mine effective feature representations and boost the segmentation performance. To produce more accurate segmentation maps, we propose a Dual-scale Complementary Fusion (DCF) module that integrates features from two scale-specific decoders operating at different scales. Extensive experimental results on five benchmark datasets demonstrate the effectiveness of our DEC-Seg against other state-of-the-art semi-supervised segmentation approaches. The implementation code will be released at https://github.com/taozh2017/DECSeg.
@@
Polyp segmentation, semi-supervised learning, scale-enhanced consistency, cross-generative consistency
@@ I Introduction
Colorectal cancer (CRC) is one of the leading causes of death worldwide, often arising from adenomatous polyps. Therefore, early detection and removal of polyps are crucial in reducing the incidence of CRC. Polyp segmentation plays a vital role in identifying and locating polyps in an early stage, enabling clinicians to diagnose and treat them effectively. Currently, various deep learning models [1, 2, 3, 4, 5] have been developed and obtained promising performance in polyp segmentation. However, these methods typically rely on a fully supervised training strategy, which requires a large number of pixel-wise annotations. Unfortunately, annotating medical images often demands highly specialized expertise, making the process time-consuming and costly. As a solution, semi-supervised learning (SSL) provides an effective approach to leverage unlabeled data to enhance the model’s performance.
@@
Recently, various semi-supervised medical image segmentation methods [6, 7, 8, 9, 10, 11, 12, 13, 14] based on deep learning have been developed. Among these methods, the consistent constraint is a widely used strategy, by enhancing that the perturbations on unlabeled data should not significantly vary their outputs or predictions. One of the most representative frameworks is the mean teacher (MT) [15], which designs a perturbation-based consistency loss between the teacher and student models on the unlabeled examples. Inspired by MT, several improved methods focus on designing different perturbation strategies to achieve SSL segmentation. For instance, an uncertainty-aware framework (UA-MT) [10] was proposed to make the student model gradually learn more reliable targets and eliminate unreliable predictions by exploiting the uncertainty information. Chen et al. [7] developed a cross-pseudo supervision strategy, which initializes two identical networks with different weights and encourages high consistency between the predictions of two different networks with different weights for the same input. Luo et al. [16] designed an uncertainty rectified pyramid consistency (URPC) scheme to enable the segmentation model can produce consistent predictions at different scales. Besides, CLCC [17] was proposed to fuse global and patch-wise information by contrastive learning and consistency constraint, which has been applied to polyp segmentation. SLCNet [18] was proposed to utilize shape information for the training by feeding pseudo-labels into the network simultaneously. Despite the achievement made, semi-supervised polyp segmentation remains challenging due to variations in the shape, size, and location of polyps. Therefore, it is crucial to develop effective and robust semi-supervised learning strategies that leverage a large amount of unlabeled data to boost the polyp segmentation performance.
@@
To this end, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised polyp Segmentation (DEC-Seg), which fully exploits multi-scale information from the labeled and unlabeled data to enhance segmentation performance. Specifically, we first propose a Cross-level Feature Aggregation (CFA) module to integrate the adjacent layer features in the encoder, and the cross-level aggregated features are further incorporated into the scale-specific decoder. Then, we present a scale-enhanced consistency strategy to encourage the consistency of segmentation maps from the same inputs with different scales. This consistency can guide our segmentation network to learn more powerful features for handling scale variations. Moreover, a Dual-scale Complementary Fusion (DCF) module is further proposed to filter the effective information from different scale features and fuse them to generate the final prediction map. Besides, we design a scale-aware perturbation consistency to impose on different scale images. Meanwhile, a cross-generative consistency is presented to constrain the original images and perturbed ones can be reconstructed by cross-segmentation maps, which can effectively harness the knowledge from unlabeled data. Finally, we conduct the comparison experiments on five benchmark datasets, and the results demonstrate the superiority of the proposed model over other state-of-the-art semi-supervised segmentation methods.
@@
  • We propose a novel semi-supervised polyp segmentation framework, which leverages scale-enhanced consistency and cross-generative consistency to exploit the relationships between labeled and unlabeled data for boosting the segmentation model.
@@
  • A dual-scale complementary fusion module is proposed to generate better predictions through the complementation and fusion of different scale information from the two scale-specific decoders.
@@
Figure 1: Architecture of our DEC-Seg framework. Labeled data are fed to the student model to achieve the segmentation task under supervised learning. Unlabeled data and the corresponding perturbed versions are passed through the student model and teacher model, which are constrained under scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency. See Sec. III for details.
The remainder of this paper is structured as follows. In Section II, we briefly review some related works. In Section III, we give the overview framework and then detail the proposed semi-supervised segmentation model. Then, we present the experimental settings, comparison results, and ablation study in Section IV. Finally, we conclude this paper in Section V.
@@ II Related works
II-A Polyp Segmentation
Recently, numerous studies have proposed different deep learning architectures and techniques to enhance the accuracy of polyp segmentation. The UNet framework, proposed by Ronneberger et al. [19], has significantly advanced the field of medical image segmentation. It incorporates skip connections between the contracting and expanding paths, enabling the network to retain fine-grained spatial information while also capturing high-level contextual features. UNet has demonstrated state-of-the-art performance across various segmentation tasks, including liver and tumor segmentation, cell segmentation, brain MRI segmentation, and polyp segmentation. To further improve the accuracy of UNet-based models, researchers have developed several variants, such as attention UNet [20], UNet++ [1], and ResUNet++ [3]. For example, Zhou et al. [1] presented UNet++ for medical image segmentation which designs a series of nested, dense skip pathways to reduce the semantic gap between the feature maps. Jha et al. [3] developed a ResUNet++ architecture, which takes advantage of residual blocks, squeeze and excitation blocks, atrous spatial pyramidal pooling, and attention blocks. Besides, several polyp segmentation methods [21, 22, 23, 24] have been developed and achieved promising performance.
@@ II-B Semi-Supervised Learning
Existing SSL methods can be categorized into self-training, consistency learning, contrastive learning, and adversarial learning. Self-training [25, 26] methods use the predictions of the fully-supervised algorithm to act as the pseudo labels of the unlabeled data and retrain the model by mixing with the labeled data. Consistency learning methods [16, 27, 15, 7] enforce agreement between different views to improve the model’s performance. By utilizing a consistency loss, these methods ensure that the model’s predictions remain consistent across different representations or augmentations of the same input. Contrastive learning methods [28, 17, 29, 30, 31] obtain better representation learning by contrasting similar features against dissimilar features. Adversarial learning methods [32, 33, 34] employ a minimax game between a generator and a discriminator to train the model. By generating new and realistic samples, these methods enhance the model’s ability to generalize and produce accurate segmentation. Several SSL methods have been developed for the medical image segmentation task. For example, Li et al. [27] proposed a segmentation model with multiple auxiliary decoders and encourage the consistency of the predictions made by the main decoder and the auxiliary decoders. Besides, contrastive learning [28, 17] has been widely used in SSL medical image segmentation. Moreover, adversarial learning [32, 33, 35] has also emerged to improve the robustness of the model. For example, Peiris et al. [32] equipped the network with a critic network to influence the segmentation network to produce the resemble prediction map as ground truth. Lei et al. [33] proposed double discriminators which are used to learn the prior relationship between labeled and unlabeled data. However, these semi-supervised methods do not deeply explore the exploitability and importance of scale information for semi-supervised learning.
@@ II-C Consistency Learning
Consistency learning as one of the most popular methods in semi-supervised learning aims to force the network to learn potential knowledge from different predictions by applying consistency constraints, which may come from different views of the same input from the same network, or may come from the same input from different networks. Currently, several methods [10, 7, 36, 37, 38] based on the mean teacher framework focus on designing different consistent and perturbation strategies to achieve SSL segmentation. For instance, Verma et al. [36] regularized semi-supervised learning by encouraging consistent predictions at interpolations of unlabeled points u1 and u2. Wu et al. [37] proposed three decoders and a different upsampling module in each decoder to amplify the perturbation and impose consistency across the three prediction maps. Zhong et al. [38] proposed a multi-attention tri-branch network (MTNet) that each branch uses a different attention mechanism. Among these methods, the consistent constraint is a widely used strategy, by enhancing that the perturbations on unlabeled data should not significantly vary their outputs.
@@
Figure 2: The detailed architectures of the encoder (i.e., $E_{s}$ and $E_{t}$ in Fig. 1) and scale-specific decoder (i.e., $D_{s1}$, $D_{s2}$, $D_{t1}$, and $D_{t2}$ in Fig. 1) networks. “Conv” denotes a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation, and “Trans-Conv” denotes a transpose convolution.
@@ III-A Overview
[…]
@@ III-B Cross-level Feature Aggregation Module
To fully exploit cross-level information and provide more effective supplementary from different convolutional layers, we propose the CFA module to fuse the features from every two adjacent layers. Then, the aggregated feature can be incorporated into our designed decoder, and the feature flow is shown in Fig. 2. Specifically, as shown in Fig. 3, two adjacent features $F_{i}$ and $F_{i+1}$ first proceeded by a $1\times 1$ convolution, and then the two features are concatenated (i.e., $F_{cat}$) to obtain $F_{cat}^{{}^{\prime}}=\mathcal{B}_{conv3\times 3}(F_{cat})$, where $\mathcal{B}_{conv3{\times}3}(\cdot)$ is a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation. To learn the cross-level attention-based enhanced feature, we conduct a global average pooling (GAP) on the cascaded feature $F_{cat}^{{}^{\prime}}$ and utilize the point-wise convolution (PWC) [40] to capture channel interactions across different spatial positions. Therefore, we obtain the attention-based weights by
@@
Figure 3: Illustrations of the encoder and cross-level feature aggregation module. “GAP” denotes the global average pooling, and “PWC” denotes the point-wise convolution.
@@ III-C Scale-enhanced Consistency
The scale variation is still a great challenge for polyp segmentation. As discussed in [41], reducing the gap between the network outputs from different scale images can help the model learn scale-correlated features. Motivated by this observation, we propose a scale-enhanced consistency scheme to constrain the outputs of different scale images to be closed, in which the different scale features can be refined by each other. Specifically, given unlabeled images $X_{u}$, $X_{u}$ and its downsampled versions $X_{u}^{d}=Down(X_{u})$ are fed to a weight-shared encoder to extract two sets of multi-level features, and then we can obtain the predicted maps $S_{u}$ and $S_{u}^{d}$, respectively. Specifically, the two segmentation maps can be obtained by $S_{u}=D_{s1}(E_{s}(X_{u}))$ and $S_{u}^{d}=D_{s2}(E_{s}(Down(X_{u})))$, where $E_{s}(\cdot)$ is the encoder in the student model, and $Down(\cdot)$ denotes a $1/2$ downsampling operation. $D_{s1}$ and $D_{s2}$ are scale-specific decoders in the student model, which are independent of each other to increase the perturbation. Similarly, for labeled images $X_{l}$, we can obtain the predicted segmentation maps $S_{l}$ and $S_{l}^{d}$ from $X_{l}$ and its downsampled versions $X_{l}^{d}$ by $S_{l}=D_{s1}(E_{s}(X_{l}))$ and $S_{l}^{d}=D_{s2}(E_{s}(Down(X_{l})))$, respectively. Therefore, we form a scale-enhanced consistency loss ($\mathcal{L}_{SC}$) to enforce $S$ and $S^{d}$ to be consistent, which is defined by
$\mathcal{L}_{SC}=\mathcal{L}_{mse}(S_{u},S_{u}^{d})+\mathcal{L}_{mse}(S_{l},S_{l}^{d}),$ (3)
where $\mathcal{L}_{mse}$ denotes the widely used mean square error loss. Finally, the supervised loss is formulated by
$\begin{split}\mathcal{L}_{S}=&\mathcal{L}_{sup}(S_{l},Y_{l})+\mathcal{L}_{sup}(S_{l}^{d},Down(Y_{l})).\end{split}$ (4)

### v2

# Dual-scale Enhanced and Cross-generative Consistency Learning for Semi-supervised Medical Image Segmentation

@@ Abstract
Medical image segmentation plays a crucial role in computer-aided diagnosis. However, existing methods heavily rely on fully supervised training, which requires a large amount of labeled data with time-consuming pixel-wise annotations. Moreover, accurately segmenting lesions poses challenges due to variations in shape, size, and location. To address these issues, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised medical image Segmentation (DEC-Seg). First, we propose a Cross-level Feature Aggregation (CFA) module that integrates cross-level adjacent layers to enhance the feature representation ability across different resolutions. To address scale variation, we present a scale-enhanced consistency constraint, which ensures consistency in the segmentation maps generated from the same input image at different scales. This constraint helps handle variations in lesion sizes and improves the robustness of the model. Furthermore, we propose a cross-generative consistency scheme, in which the original and perturbed images can be reconstructed using cross-segmentation maps. This consistency constraint allows us to mine effective feature representations and boost the segmentation performance. To further exploit the scale information, we propose a Dual-scale Complementary Fusion (DCF) module that integrates features from two scale-specific decoders operating at different scales to help produce more accurate segmentation maps. Extensive experimental results on multiple medical segmentation tasks (polyp, skin lesion, and brain glioma) demonstrate the effectiveness of our DEC-Seg against other state-of-the-art semi-supervised segmentation approaches. The implementation code will be released at https://github.com/taozh2017/DECSeg.
@@
Medical image segmentation, semi-supervised learning, scale-enhanced consistency, cross-generative consistency.
@@ I Introduction
Accurate segmentation of organs or lesions from medical images is a crucial task within the realm of medical image processing and analysis. The successful accomplishment of this task holds immense significance for disease diagnosis, organ localization, and lesion segmentation. In recent years, several deep learning models [1, 2, 3, 4] have been developed and have shown promising performance in medical image segmentation. However, these models often rely on a fully supervised training strategy, which necessitates a substantial number of pixel-wise annotations. Unfortunately, the annotation process for medical images is both time-consuming and costly due to the need for specialized expertise. To address these challenges, researchers have turned to semi-supervised learning (SSL) as an effective approach to leverage unlabeled data and enhance the performance of segmentation models [5, 6]. By utilizing SSL techniques, models can exploit the abundance of available unlabeled data in medical imaging datasets, reducing their dependence on fully annotated data. This approach not only alleviates the burden of extensive annotation requirements but also has the potential to improve the model’s accuracy in medical image segmentation tasks.
@@
Recently, various semi-supervised medical image segmentation methods [7, 8, 9] based on deep learning have been developed. Among these methods, the consistent constraint is a widely used strategy, by ensuring that the perturbations on unlabeled data should not significantly vary their outputs or predictions. One of the most representative frameworks is the mean teacher (MT) [10], which designs a perturbation-based consistency loss between the teacher and student models on the unlabeled examples. Inspired by MT, several improved methods focus on designing different perturbation strategies to achieve SSL segmentation. For instance, an uncertainty-aware framework (UA-MT) [11] was proposed to make the student model gradually learn more reliable targets and eliminate unreliable predictions by exploiting the uncertainty information. Verma et al. [12] proposed an interpolation consistency training framework, which constrains the prediction at an interpolation of unlabeled points to be consistent with that of the predictions at those points. Chen et al. [13] developed a cross-pseudo supervision strategy, which initializes two identical networks with different weights and encourages high consistency between the predictions of two different networks with different weights for the same input. Luo et al. [14] designed an uncertainty rectified pyramid consistency (URPC) scheme to enable the segmentation model can produce consistent predictions at different scales. Besides, CLCC [15] was proposed to fuse global and patch-wise information by contrastive learning and consistency constraint. SLCNet [16] was proposed to utilize shape information for the training by feeding pseudo-labels into the network simultaneously. Despite the achievement made, semi-supervised medical image segmentation remains challenging due to variations in the shape, size, and location of lesions. Therefore, it is crucial to develop effective and robust semi-supervised learning strategies that leverage a large amount of unlabeled data to boost the medical image segmentation performance.
@@
To this end, we propose a novel Dual-scale Enhanced and Cross-generative consistency learning framework for semi-supervised medical image Segmentation (DEC-Seg), which fully exploits multi-scale information from the labeled and unlabeled data to enhance segmentation performance. Specifically, we first propose a Cross-level Feature Aggregation (CFA) module to integrate the adjacent layer features in the encoder, and the cross-level aggregated features are further incorporated into the scale-specific decoder. Then, we present a scale-enhanced consistency strategy to encourage the consistency of segmentation maps from the same inputs with different scales. This consistency can guide our segmentation network to learn more powerful features for handling scale variations. Moreover, a Dual-scale Complementary Fusion (DCF) module is further proposed to filter the effective information from different scale features and fuse them to boost segmentation performance. Meanwhile, a cross-generative consistency is presented to constrain the original images and perturbed ones can be reconstructed by cross-segmentation maps, which can effectively harness the knowledge from unlabeled data. Finally, we conduct the comparison experiments on three medical image segmentation tasks, and the results demonstrate the superiority of the proposed model over other state-of-the-art semi-supervised segmentation methods.
@@
  • We propose a novel semi-supervised medical image segmentation framework, which leverages scale-enhanced consistency and cross-generative consistency to exploit the relationships between labeled and unlabeled data for boosting the segmentation model.
@@
  • A dual-scale complementary fusion module is proposed to benifit the better prediction generation through the complementation and fusion of different scale information from the two scale-specific decoders.
@@
Figure 1: Architecture of our DEC-Seg framework. Labeled data are fed to the model to achieve the segmentation task under supervised learning. Unlabeled data and the corresponding perturbed versions are passed through the model, which is constrained under scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency. See Sec. III for details.
The remainder of this paper is structured as follows. In Section II, we briefly review some related works to our model, consisting of medical image segmentation, semi-supervised learning, and consistency learning. In Section III, we give the overview framework and then provide the details of the proposed semi-supervised segmentation model. Then, we present the used datasets, experimental settings, comparison results, and ablation study in Section IV. Finally, we conclude this paper in Section V.
@@ II Related Works
II-A Medical Image Segmentation
Recently, numerous studies have proposed different deep learning architectures and techniques to enhance the accuracy of medical image segmentation. The UNet framework, proposed by Ronneberger et al. [17], has significantly advanced the field of medical image segmentation. It incorporates skip connections between the contracting and expanding paths, enabling the network to retain fine-grained spatial information while also capturing high-level contextual features. UNet has demonstrated state-of-the-art performance across various segmentation tasks, including liver and tumor segmentation, cell segmentation, brain tumor segmentation, and polyp segmentation. To further improve the accuracy of UNet-based models, researchers have developed several variants, such as attention UNet [18], UNet++ [1], and ResUNet++ [19]. For example, Zhou et al. [1] presented UNet++ for medical image segmentation which designs a series of nested, dense skip pathways to reduce the semantic gap between the feature maps. Jha et al. [19] developed a ResUNet++ architecture, which takes advantage of residual blocks, squeeze and excitation blocks, atrous spatial pyramidal pooling, and attention blocks. Moreover, the cross-level and multi-level feature fusion technique is widely recognized for enhancing segmentation performance by integrating spatial information from low-level features with semantic information from high-level features. For instance, Li et al. [20] proposed a cross-level information processing module that aggregates and processes multi-scale features transmitted by different encoder layers to improve polyp segmentation. Zhou et al. [21] designed a cross-level feature fusion module to merge adjacent features from various levels, enabling the characterization of cross-level and multi-scale information to address scale variations in polyps. Additionally, CFU-Net [22] incorporates a multi-level attention module to exploit interactions of contextual information, decision information, and long-range dependencies, thereby enhancing feature propagation within skip connections.
@@ II-B Semi-Supervised Learning
Existing SSL methods can be categorized into self-training, consistency learning, contrastive learning, and adversarial learning. Self-training [23, 24] methods use the predictions of the fully-supervised algorithm to act as the pseudo labels of the unlabeled data and retrain the model by mixing with the labeled data. Consistency learning methods [14, 25, 10, 13] enforce agreement between different views to improve the model’s performance. By utilizing a consistency loss, these methods ensure that the model’s predictions remain consistent across different representations or augmentations of the same input. Contrastive learning methods [26, 15, 27, 28, 29] obtain better representation learning by contrasting similar features against dissimilar features. Adversarial learning methods [30, 31, 32] employ a minimax game between a generator and a discriminator to train the model. By generating new and realistic samples, these methods enhance the model’s ability to generalize and produce accurate segmentation. Several SSL methods have been developed for the medical image segmentation task. For example, Li et al. [25] proposed a segmentation model with multiple auxiliary decoders and encouraged the consistency of the predictions made by the main decoder and the auxiliary decoders. Besides, contrastive learning [26, 33, 34] has been widely used in SSL medical image segmentation. Moreover, adversarial learning [30, 31, 35] has also emerged to improve the robustness of the model. For example, Peiris et al. [30] equipped the network with a critic network to influence the segmentation network to produce the resemble prediction map as ground truth. Lei et al. [31] proposed double discriminators which are used to learn the prior relationship between labeled and unlabeled data. However, these semi-supervised methods do not deeply explore the exploitability and importance of scale information for semi-supervised learning.
@@ II-C Consistency Learning
Consistency learning as one of the most popular methods in semi-supervised learning aims to force the network to learn potential knowledge from different predictions by applying consistency constraints, which may come from different views of the same input from the same network, or may come from the same input from different networks. Currently, several methods [10, 13, 12, 36, 37] based on the mean teacher framework focus on designing different consistent and perturbation strategies to achieve SSL segmentation. For instance, Tarvainen et al. [10] proposed to generate two different augmented views of the same input and use the output of the teacher model to supervise the student model. Chen et al. [13] proposed cross-pseudo supervision by enforcing the pseudo supervision between the predictions of two models with different parameter initialization. Wu et al. [36] proposed three decoders and a different upsampling module in each decoder to amplify the perturbation and impose consistency across the three prediction maps. Zhong et al. [37] proposed a multi-attention tri-branch network (MTNet) that each branch uses a different attention mechanism. Among these methods, the consistent constraint is a widely used strategy, aiming to ensure that perturbations applied to unlabeled data do not drastically alter their outputs.
@@
Figure 2: The detailed architectures of the encoder (i.e., $E$ in Fig. 1) and scale-specific decoder (i.e., $D_{1}$ and $D_{2}$ in Fig. 1) networks. “Conv” denotes a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation, and “Trans-Conv” denotes a transpose convolution.
@@ III-A Overview
The proposed DEC-Seg framework, as depicted in Fig. 1, fully leverages multi-scale information from both labeled and unlabeled data for medical image segmentation. Our segmentation network comprises a shared encoder ($E$), two scale-specific decoders ($D_{1}$ processes features solely from the original scale while $D_{2}$ exclusively processes features from the downsampled scale), and a scale-fused decoder ($D_{f}$ is used to fuse features from both scales). Additionally, two generative networks ($G_{1}$ and $G_{2}$) assist in the segmentation process, enhancing the feature mining ability of our segmentation network through a cross-generation strategy. Specifically, the original and downsampled images from the labeled data are fed into the model (i.e., the encoder $E$ and scale-specific decoders $D_{1}$ and $D_{2}$), and then the scale-enhanced consistency and supervised loss are calculated to constrain the network. Then, the unlabeled data and its perturbed versions are inputted into the model. Meanwhile, we conduct scale-enhanced consistency, scale-aware perturbation consistency, and cross-generative consistency for unlabeled data, to fully leverage unlabeled samples to improve the segmentation performance. Moreover, to make use of multi-scale information, we design the scale-fused decoder (i.e., Df) to generate the final segmentation maps. For convenience, we denote ${X}_{l}=\left\{x_{l}^{1},x_{l}^{2},\cdots,x_{l}^{n_{l}}\right\}$ and ${Y}_{l}=\left\{y_{l}^{1},y_{l}^{2},\cdots,y_{l}^{n_{l}}\right\}$ as the labeled dataset and the corresponding label set, respectively, where $n_{l}$ is the number of the labeled images. We use ${X}_{u}=\left\{x_{u}^{1},x_{u}^{2},\cdots,x_{u}^{n_{u}}\right\}$ to denote the unlabeled dataset with $n_{u}$ images, and we typically have $n_{l}\ll n_{u}$. Given inputs $X$, they are fed into the encoder to learn five levels of features, namely $\{F_{i}\}_{i=1}^{5}$.
@@ III-B Cross-level Feature Aggregation Module
To fully exploit cross-level information from various receptive fields and provide more effective supplementation from different convolutional layers, we propose the CFA module. This module fuses features from every two adjacent layers, enhancing the understanding of context and relationships within the cross-level features. The aggregated feature is then incorporated into our designed decoder, as depicted in Fig. 2. Specifically, as shown in Fig. 3, two adjacent features $F_{i}$ and $F_{i+1}$ first proceeded by a $1\times 1$ convolution, and then the two features are concatenated (i.e., $F_{cat}$) to obtain $F_{cat}^{{}^{\prime}}=\mathcal{B}_{conv3\times 3}(F_{cat})$, where $\mathcal{B}_{conv3{\times}3}(\cdot)$ is a sequential operation that consists of a $3\times{3}$ convolution, batch normalization, and a ReLU activation. To learn the cross-level attention-based enhanced feature, we conduct a global average pooling (GAP) on the cascaded feature $F_{cat}^{{}^{\prime}}$ and utilize the point-wise convolution (PWC) [38] to capture channel interactions across different spatial positions. Therefore, we obtain the attention-based weights by
@@
Figure 3: Illustrations of cross-level feature aggregation module. “GAP” denotes the global average pooling, “PWC” denotes the point-wise convolution, and “$U$” denotes the upsampling using bilinear interpolation.
@@ III-C Scale-enhanced Consistency
The scale variation is still a great challenge for medical image segmentation. As discussed in [39], reducing the gap between the network outputs from different scale images can help the model learn scale-correlated features. Motivated by this observation, we propose a scale-enhanced consistency scheme to constrain the outputs of different scale images to be closed, in which the different scale features can be refined by each other. To increase the perturbation, we use two independent decoders for the single-scale prediction task, which are called scale-special decoders (i.e., $D_{1}$ and $D_{2}$). Specifically, given unlabeled images $X_{u}$, $X_{u}$ and its downsampled versions $X_{u}^{2}=Down(X_{u})$ are fed to a weight-shared encoder to extract two sets of multi-level features, and then we can obtain the predicted maps $S_{u}^{1}$ and $S_{u}^{2}$, respectively. Specifically, the two segmentation maps can be obtained by $S_{u}^{1}=D_{1}(E(X_{u}))$ and $S_{u}^{2}=D_{2}(E(X_{u}^{2}))$, where $E(\cdot)$ is the encoder, and $Down(\cdot)$ denotes a $1/2$ downsampling operation. Furthermore, we employ the mutual supervision approach outlined in [13] to ensure the consistency of $S^{1}_{u}$ and $S^{2}_{u}$ with different sizes. Concretely, $\hat{Y}^{1}_{u}=\arg\max\limits_{c}(S^{1}_{u})$ and $\hat{Y}^{2}_{u}=\arg\max\limits_{c}(S^{2}_{u})$ are computed. Subsequently, $\hat{Y}^{1}_{u}$ is downsampled to the size of $S^{2}_{u}$ and utilized as a pseudo-label for $S^{2}_{u}$, while $\hat{Y}^{2}_{u}$ is upsampled to the size of $S^{1}_{u}$ and employed as a pseudo-label for $S^{1}_{u}$. Therefore, the scale-enhanced consistency loss for unlabeled data is defined by
$\mathcal{L}_{SC}^{u}(S^{1}_{u},S^{2}_{u})=\mathcal{L}_{ce}(S^{1}_{u},Up(\hat{Y}^{2}_{u}))+\mathcal{L}_{ce}(S^{2}_{u},Down(\hat{Y}^{1}_{u})),$ (3)
where $\mathcal{L}_{ce}$ denotes the widely used cross entropy loss and $Up(\cdot)$ denotes a 2×2$\times$2 × upsampling operation. Similarly, for labeled images $X_{l}$, we can obtain the predicted segmentation maps $S_{l}^{1}$ and $S_{l}^{2}$ from $X_{l}$ and its downsampled versions $X_{l}^{2}$ by $S_{l}^{1}=D_{1}(E(X_{l}))$ and $S_{l}^{2}=D_{2}(E(X_{l}^{2}))$, respectively. As a result, $\mathcal{L}_{SC}^{l}(S^{1}_{l},S^{2}_{l})$ can be obtained in the same way.
[…]
Figure 4: Illustration of (a) the architecture of the scale-fused decoder (i.e., ${D}_{f}$ in Fig. 1) and (b) the proposed dual-scale complementary fusion module. “$U$” denotes the upsampling using a bilinear interpolation.